OpenStudy (anonymous):

what equals 7 3/5 x 1 3/7?

OpenStudy (anonymous):

can you make the mixed number into a improper farcion

OpenStudy (anonymous):

if you want to make it into an improper fraction you must first do this \[7*5 + 3=38/5\]\[1*7+3=10/7\] then 38/5 * 10/7 = 380/35 = 76/5